OpenStudy (anonymous):

When using FOIL to multiply the binomials (2x – 1)(–3x + 4), what is the value of the "inner" terms?

OpenStudy (zzr0ck3r):

\((a+b)(c+d)=ac+ad+bc+bd\)

OpenStudy (unklerhaukus):

\[(\overbrace{\overbrace{\color{orange}a +\underbrace{\color{cornflowerblue}b)(\color{seagreen}c}_{\text{Inside}}}^{\text{First}} +\color{brown}d}^{\text{Outside}}) =\overbrace{\color{orange}a\color{seagreen}c}^\text F +\overbrace{\color{orange}a\color{brown}d}^\text O +\underbrace{\color{cornflowerblue}b\color{seagreen}c}_\text I +\underbrace{\color{cornflowerblue}b\color{brown}d}_\text L\\ \qquad\quad \underbrace{\qquad\qquad}_{\text{Last}}\]

OpenStudy (zzr0ck3r):

really its just distribution \((a+b)(c+d)=(a+b)c+(a+b)d=ac+bc+ad+bd\)

OpenStudy (anonymous):

(2x)(-3x+4) + (-1)(-3x+4)

OpenStudy (zzr0ck3r):

holly god that is pretty @UnkleRhaukus

OpenStudy (zzr0ck3r):

I cant see others typing on my tablet or I would not have said anything

OpenStudy (anonymous):

So ok I am a little confused here >.> its a multiple choice answer A. –4 B. –4x C. –3x D. 3x would it be B?

OpenStudy (zzr0ck3r):

wow its all lined up so its in order F O I L @UnkleRhaukus

OpenStudy (zzr0ck3r):

you get my response of the year award

Miracrown (miracrown):

Here we would have 3x or (D).

OpenStudy (anonymous):

Thank you!!! everyone for your help!!

Miracrown (miracrown):

Multiplying together these 2 terms gives us 3x.

Miracrown (miracrown):